### v2

# Adapting Pre-Trained Vision Models for Novel Instance Detection and Segmentation

Abstract
Novel Instance Detection and Segmentation (NIDS) aims at detecting and segmenting novel object instances given a few examples of each instance. We propose a unified, simple yet effective framework (NIDS-Net) comprising object proposal generation, embedding creation for both instance templates and proposal regions, and embedding matching for instance label assignment. Leveraging recent advancements in large vision methods, we utilize Grounding DINO and Segment Anything Model (SAM) to obtain object proposals with accurate bounding boxes and masks. Central to our approach is the generation of high-quality instance embeddings. We utilize foreground feature averages of patch embeddings from the DINOv2 ViT backbone, followed by refinement through a weight adapter mechanism that we introduce. We show experimentally that our weight adapter can adjust the embeddings locally within their feature space and effectively limit overfitting in the few-shot setting. This methodology enables a straightforward matching strategy, resulting in significant performance gains. Our framework surpasses current state-of-the-art methods, demonstrating notable improvements of 22.3, 46.2, 10.3, and 24.0 in average precision (AP) across four detection datasets. In instance segmentation tasks on seven core datasets of the BOP challenge, our method is around 4.5 times faster than the leading published RGB method and surpasses it by 3.6 AP. NIDS-Net is about 5.7 times faster than the top RGB-D method while maintaining competitive performance [^1] Project Page: https://irvlutd.github.io/NIDSNet/.
1 Introduction
Figure 1: We leverage pre-trained vision models for object proposal generation and feature extraction, and introduce a weight adapter to improve pre-trained feature embeddings for novel object instance detection and segmentation.
Novel Instance Detection and Segmentation (NIDS) is a crucial task in computer vision aimed at identifying and locating unseen instances in images or videos, given a few examples of each instance. It also has wide applications in robotics. For example, suppose a robot needs to grasp a specific, novel object instance from a cluttered desk given only a small number of multi-view template images of the object. Ideally, detectors for this task should provide precise bounding boxes and masks of the target instances given a query image. However, standard detection methods trained on specific classes often struggle to recognize new objects.
The current paradigm for NIDS typically encompasses the following steps: (1) generating proposals from a query image, (2) obtaining embeddings of the proposals and the object templates, and (3) matching the embeddings of proposals with those of the templates for identification. Recent work [49, 81, 66, 53] has utilized various open-world detectors to obtain object proposals. However, the use of models such as the Segment Anything Model (SAM) [47] or FastSAM [103] often results in the generation of region-based proposals rather than actual object proposals. This misidentification of regions as objects can lead to object identification issues. For example, a single object may be divided into multiple proposals or some background regions may be mistaken for foreground objects. These false alarms not only slow down the whole detection progress but also spoil the final results.
To embed proposals and templates, some existing works, e.g., [81, 66, 53], adopt the $cls$ token of DINOv2  [68, 20] while others, e.g., Li et al. [49], employ a 3D voxel representation. Ideally, for each specific unseen instance, the template embeddings from different views should be similar to each other but markedly different from the embeddings of other instances. Despite these efforts, the embeddings of different instances may remain similar, and so the challenges persist.
To address the limitations of the existing methods, we propose a framework for NIDS, as depicted in Fig. 1. Initially, we utilize Grounding DINO [56] with the text prompt “objects” on a cluttered query image to obtain bounding boxes for foreground objects. This method capitalizes on the inherent objectness of Grounding DINO using specialized prompts rather than standard labels. Subsequently, we employ SAM to create masks within these bounding boxes, generating precise object proposals comprising both bounding boxes and masks. For instance embeddings, we first extract average foreground features [48] from the patch embeddings of the DINOv2 ViT backbone. We then apply adapters to enhance instance embeddings by clustering similar instances and distancing different ones.
The adapters are trained using the template images only, where InfoNCE loss [67, 15] is applied to the feature embeddings after the adapters. To refine these embeddings, the CLIP-Adapter [29] introduces a residual vector added to the original embedding, which risks overfitting with a few training examples. This addition can spoil and destabilize the embeddings of background objects, disrupting the entire feature space. To mitigate this issue, we, instead, introduce a weight adapter that modifies the original embeddings by applying learned weights. Finally, we employ a straightforward matching approach [62], such as stable matching or using $argmax$, to assign instance labels to these proposals.
Our approach has been validated on four detection datasets and seven segmentation datasets for NIDS. The framework significantly surpasses existing state-of-the-art methods: it demonstrates substantial increases in average precision (AP), with gains of 22.3, 46.2, 10.3, and 24.0 across these detection datasets. Moreover, in instance segmentation tasks on the seven core datasets of the BOP challenge [84], our method outperforms the best published RGB and RGB-D methods.
Our contributions are summarized as follows.
  • We propose NIDS-Net, a unified framework for novel instance detection and segmentation that surpasses existing published methods across four detection datasets and seven segmentation datasets of BOP challenge.
  • We leverage the objectness of the open-world detector Grounding DINO and region recognition of Segmentation Anything Model (SAM) to obtain object proposals.
  • We introduce the Weight Adapter, a tool designed to refine embeddings within their feature space while discouraging overfitting. This approach yields increasingly significant improvements as the adapting of the feature space becomes more robust and stable.
2 Related Work
Pretrained Models. Large-scale pretraining has demonstrated utility across various downstream tasks. Works such as [32, 91, 59, 24, 58, 10, 68] emphasize large-scale image representation learning, while [74, 50] focus on vision-language pretraining. Additionally, [3, 4, 96] address depth prediction, and [72, 73, 6, 87] tackle large-scale language modeling. Object detection is addressed by [79, 77, 43, 44, 31, 56], while image segmentation is the focus of [33, 89, 46]. Works such as [94, 52, 14, 76] explore text-to-3D generation, and [82, 1] address text-to-video synthesis. Finally, 6DoF pose estimation and tracking are tackled by [90, 21, 51, 88]. DINOv2 [68, 20] offers robust features to represent unseen instances. These foundational models offer diverse capabilities, with the challenge lying in effectively leveraging their wealth of knowledge for specific use cases, such as novel instance detection. In our work, we leverage pre-trained vision models for NIDS.
Open-World Object Detection (OWOD). OWOD tackles real-world scenarios where object detection algorithms are tasked not only with identifying known and labeled objects but also with effectively handling novel and unknown objects during inference. This sets OWOD apart from traditional object detection, which confines itself to detecting only predetermined object classes. A key challenge in OWOD lies in detecting and classifying unknown objects that were not part of the training data.
Recent approaches such as OWL-ViT [65], DITO [43], RO-ViT [44], YOLO-World [18], Sambor [31], and Grounding DINO [56] offer diverse solutions to address OWOD challenges. For example, OWL-ViT leverages a standard Vision Transformer [24] architecture with minimal modifications, incorporating contrastive image-text pre-training and end-to-end detection fine-tuning. DITO introduces an open-vocabulary detection approach based on detection-oriented image-text pretraining, replacing the classification architecture with a detector architecture during pretraining to better serve region-level recognition needs. Grounding DINO [56] utilizes features from DINO [10] to ground text queries. We employ Grounding DINO due to its strong performance. Similar to our work, Fang et al. [28] obtain object bounding boxes using the text prompt “a rigid object” with Grounding DINO.
Few-Shot Object Detection (FSOD). FSOD involves detecting occurrences of a novel class, unseen during training, given only a few template examples. For example, De-ViT [102] introduces a region-propagation mechanism for localization, transforming propagated region masks into bounding boxes using a learnable spatial integral layer. FS-DETR [7], the first few-shot detection transformer, extends DETR [9] to process arbitrary numbers of novel objects concurrently while supporting various examples from each class. Other approaches such as [27] prioritize salient representations through angle distance and magnitude dispersion across few-shot samples,  [93] leverages variational autoencoders for data augmentation, and  [61] learns geometry-aware features with an offline simplex equiangular tight frame classifier. Additionally,  [60] generates specific and representative prototypes for each query image, considering both support and query image features. Our method is a FSOD method based on region proposal generation and matching. Compared to previous FSOD methods, it is simple yet effective.
Instance Detection. Instance detection involves identifying an unseen instance in a test image using corresponding templates. Some methods, e.g.,  [64, 2, 57], rely on pure 2D representations and matching techniques. Specifically, DTOID [64] employs global object attention and a local pose-specific branch for template-guided heatmap prediction. However, these methods may struggle with variations in 2D appearance caused by occlusion or pose changes. In contrast, VoxDet [49] utilizes explicit 3D knowledge from multi-view templates, providing geometry-invariant representations. We generate 2D robust instance embeddings from templates with DINOv2 [68, 20].
Adapters for Pre-trained Models. The application of adapters atop large pre-trained models has emerged as a prominent strategy, yielding significant improvements across various tasks. Previous works have leveraged adapters to enhance few-shot image classification tasks [29, 99, 70, 101, 83]. Other works have demonstrated effectiveness in language model fine-tuning [37, 39, 100, 30]. Additional works exhibit the promise of adapters in vision tasks including detection and segmentation [17, 16, 71, 92, 23]. For novel instance detection and segmentation, we propose the Weight Adapter to enhance performance. The adapter assigns weights to the embeddings from frozen backbones. SENets [38] employ the Squeeze-and-Excitation (SE) blocks to distribute weights. In contrast, our adapter is an independent model with a structural difference from the SE block. For instance, the SE block lacks a ReLU layer prior to the sigmoid function.
3 Method
Novel Instance Detection and Segmentation (NIDS), also referred to as open-world detection and segmentation, is a task to locate and label unseen instances within a query image. We will assume that each of the $N$ target instances is represented by $K$ template images $I_{T}\in\mathbb{R}^{K\times 3\times W\times H}$ and their corresponding segmentation masks. During inference, the output of each query image $I_{Q}\in\mathbb{R}^{3\times W\times H}$ provides the bounding boxes for detection and instance masks for segmentation of these instances.
3.1 Instance Embedding Generation Stage
In our approach, an instance embedding summarizes the pixels of an instance. The objective of this stage is to generate initial instance embeddings. Each of the $N$ instances has $K$ multi-view template images and their ground truth segmentation masks from which we derive template embeddings $E_{T}\in\mathbb{R}^{N\times K\times C}$, where $C$ is the dimension of the embeddings. Given an image and its corresponding mask, we initially extract patch embeddings using the ViT backbone of DINOv2 [68], and subsequently obtain foreground features as specified by the mask. We then perform average pooling on these features. This process, termed Foreground Feature Averaging (FFA), is proposed by [48] to assess object similarity. We employ FFA to generate all initial instance embeddings using the object templates.
3.2 Object Proposal Stage
The purpose of this stage is to acquire object proposals from a query image. Previous studies [81, 66, 53] have utilized SAM [47] or FastSAM [103] to generate regions as object proposals. Such proposals contain a high number of false alarms. For example, SAM might misclassify background regions or parts of objects as complete objects. To address this challenge, an off-the-shelf zero-shot detector, Grounding Dino [56], is employed with the text prompt “objects” to obtain initial bounding boxes of foreground objects. Then, SAM is applied to create masks based on these bounding boxes. The integration of these two models, termed Grounded-SAM (GS) [80], significantly reduces the number of erroneous object proposals and expedites the subsequent stages. Moreover, this approach eliminates the need for training of the detector. We then extract the proposal regions along with their corresponding masks from the query images. Using the FFA pipeline with proposals, we calculate proposal embeddings $E_{P}\in\mathbb{R}^{Q\times C}$, which correspond to $Q$ regions of interest. Fig. 2 illustrates the process of obtaining proposal embeddings.
Figure 2: In our framework NIDS-Net, only adapters are learnable, while other models are frozen. Instance IDs are the instance labels.
3.3 Embedding Refinement via an Adapter
Most instance embeddings are well separated. However, some embeddings from distinct instances may cluster together. For example, background object embeddings might be similar to targets. Some target instances may resemble each other. To address this issue, we employ learnable adapters to refine the embeddings. We train these adapters using the InfoNCE loss [67, 15], aiming to bring the embeddings of the same instance closer together while separating the embeddings of different instances. This training only uses the template images.
CLIP-Adapter (CA). CLIP-Adapter [29] comprises two trainable linear bottleneck layers appended to the language and image branches of CLIP. During few-shot fine-tuning, the original CLIP backbone remains frozen. Nonetheless, adding extra layers for fine-tuning may result in overfitting to the few examples available. To mitigate this issue, CLIP-Adapter integrates residual connections that dynamically merge the newly adapted knowledge with the foundational knowledge from the original CLIP backbone.
Given an instance embedding $f$ and an $MLP$ (Multi-Layer Perceptron), the CLIP-Adapter modifies the embedding according to the formula $f_{c}=\alpha\times MLP(f)+(1-\alpha)\times f$. In this equation, $f_{c}$ represents the adapted embedding, which is a linear combination of the transformed embedding $MLP(f)$ and the original embedding $f$. The residual ratio $\alpha$ is set as 0.6.
Our Weight Adapter (WA). The CLIP-Adapter refines instance embeddings by adding a new feature vector. However, this new vector might not align with the feature space of original instance embeddings, which can lead to overfitting. Since only several hundred template embeddings of target instances are trained and refined in some datasets, the embeddings of background objects can be spoiled as the process disrupts the distances between target and background instance embeddings. Assuming the feature space of original embeddings is robust and effective, it is crucial to fine-tune the instance embeddings within this space by constraining the scope of the adapted embeddings. This refinement can be accomplished by simply applying weights to the original instance embeddings.
Hence, we introduce our Weight Adapter, a compact MLP network depicted in Fig. 3. The operation of this adapter is described through the following equations:
$w$ $=\text{sigmoid}(MLP(\beta f)),$ (1)
$f_{w}$ $=w\odot(\beta f).$ (2)
Here, $\beta$ is a positive scalar. We set $\beta$ as 10. $w$ is the weights computed by passing the scaled embedding $\beta f$ through the MLP and then applying a $sigmoid$ function. The output $f_{w}$ is the element-wise product of $w$ and the scaled embedding $\beta f$, resulting in the weighted embedding.
We include a ReLU layer before the $sigmoid$ function, ensuring that the sigmoid’s output is constrained to the range $[0.5,1)$. Subsequently, each refined embedding is close to its original counterpart. Additionally, $\beta$ allows for broader adjustments within the feature space. Given that cosine similarity is utilized to compare embeddings in the subsequent phase, our adapter alters the interrelationships among the values within each embedding. Further adapter details are presented in Fig. 3.
Figure 3: Comparison of two adapters. (left) CLIP-Adapter. (right) Our weight adapter.
3.4 Matching Stage
This stage provides each proposal $i$ with an instance ID $o_{i}$ as its label and a confidence score $s_{i}$.
Initially, $Q$ proposal embeddings $E_{P}$ are matched with $N\times K$ template embeddings $E_{T}$ using cosine similarity. It yields a matrix of template scores with dimensions $Q\times N\times K$, as illustrated in Fig. 2. For each instance, we aggregate all $K$ template scores to derive a matrix of instance scores with dimensions $Q\times N$. We employ $Max$ as the aggregation function for optimal results.
Bonus Instance Score for Segmentation. To improve segmentation performance, we incorporate an additional appearance matching score, $s_{\text{appe}}$, as proposed by SAM-6D [53], into the instance scores. The final instance scores are computed as the average of $s_{\text{appe}}$ and the initial instance scores. $s_{\text{appe}}$ is used to identify objects that are semantically similar yet differ in appearance. For each proposal, we can identify its most similar template $T_{\text{best}}$ according to template scores. $s_{\text{appe}}$ is derived from the patch embeddings of a proposal image $I$ and $T_{\text{best}}$. It quantifies the similarity between the query image and the best template in terms of their local features ($f_{I,j}^{\text{patch}}$, $f_{T_{\text{best}},i}^{\text{patch}}$) as follows:
$s_{\text{appe}}=\frac{1}{N_{I}^{\text{patch}}}\sum_{j=1}^{N_{I}^{\text{patch}}}\max_{i=1,\ldots,N_{T_{\text{best}}}^{\text{patch}}}\left(\frac{f_{I,j}^{\text{patch}}\cdot f_{T_{\text{best}},i}^{\text{patch}}}{\|f_{I,j}^{\text{patch}}\|_{2}\|f_{T_{\text{best}},i}^{\text{patch}}\|_{2}}\right),$ (3)
where $N^{\text{patch}}$ is the number of patch embeddings.
Assuming that objects are unique within a query image [81], we employ the stable matching algorithm [62] on instance scores to assign a unique instance ID to each proposal. If the assumption of uniqueness is not met, we use the $Argmax$ function on the instance score matrix, which permits multiple proposals to share the same instance ID.
After matching, we acquire labeled proposals. Each is defined by $\{b_{i},M_{i},o_{i},s_{i}\}$. Here, $b_{i}$ is the bounding box of an instance $M_{i}$ is the modal mask, which covers the visible object surface [84], $o_{i}$ is the instance ID, and $s_{i}$ is the confidence score. A confidence score threshold $\delta$ can be used to remove incorrect proposals.
Table 1: Detection results on the high-resolution dataset. avg indicates that we are evaluating all images, including those from the easy and hard scenes. We train these two adapters with 2400 template embeddings.
Method AP AP50 AP75 Time (sec)
avg hard easy small medium large
FasterRCNN [78] 19.5 10.3 23.8 5.0 22.2 38.0 29.2 23.3 0.00399
RetinaNet [55] 22.2 14.9 26.5 5.5 25.8 42.7 31.2 25.0 0.00412
CenterNet [104] 21.1 11.8 25.7 5.9 24.1 40.4 32.7 23.6 0.00376
FCOS [85] 22.4 13.2 28.7 6.2 26.5 38.1 32.8 25.5 0.00271
DINO [98] 28.0 17.9 32.6 11.5 30.7 45.1 32.2 32.2 1.90625
SAM + DINOf [11] 37.0 22.4 43.9 11.9 40.9 62.7 44.1 40.4 15.10
SAM + DINOv2f [68] 41.6 28.0 47.6 14.6 45.8 69.1 49.1 46.0 14.70
NIDS-Net (Ours) 59.3 41.7 67.4 12.7 58.1 78.8 71.1 65.1 6.92
NIDS-Net + CA (Ours) 61.3 42.5 69.6 17.0 59.3 81.6 73.2 67.0 6.99
NIDS-Net + WA (Ours) 63.9 43.4 72.7 18.1 62.5 84.0 76.6 70.6 6.78
4 Experiments
We employ the Grounding DINO model with a Swin-T backbone [56] and a text prompt “objects”, and the default ViT-H SAM [47] to generate object proposals. According to the proposals, we obtain the regions of interest (RoIs) and resize them to $448\times 448$ or $224\times 224$ resolutions. For object detection, instance embeddings are produced using the DINOv2’s ViT-L model with registers [20], employing stable matching due to the uniqueness of instances in these datasets. For image segmentation, following SAM-6D [53], we utilize the ViT-L model of DINOv2 [68]. In scenarios with identical instances, such as the cluttered scenes in BOP datasets, we apply the $argmax$ function for matching.
The experiments are conducted on an RTX A5000 GPU. For each dataset, we train both adapters on template embeddings using InfoNCE loss [67, 15] with temperature 0.05. With two linear layers, both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$. The Weight Adapter is trained with the Adam optimizer [45] at a learning rate of 1e-3 and a batch size of 1024, while the CLIP-Adapter is trained at a learning rate of 1e-4. Dropout [34] is incorporated in the CLIP-Adapter to reduce overfitting. For additional details, please see the appendix.
4.1 Detection Datasets
We utilize recently developed instance detection datasets and their associated baselines [81, 49]. The High-resolution and RoboTools datasets employ real template images, while the LM-O and YCB-V datasets utilize synthetic images.
High-resolution Dataset. Shen et al. [81] design a high-resolution dataset for instance detection. This dataset comprises 100 distinct object instances, each represented by 24 photos taken from multiple viewpoints, with each photo having a resolution of $3072\times 3072$ pixels. These instances are integrated into 14 different indoor scenes for testing, captured in even higher resolution ($6144\times 8192$ pixels). The testing images are further classified as easy or hard based on the level of scene clutter and how much the instances are occluded. Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
According to Shen et al. [81], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method. Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds. This enables training detectors by viewing each instance as a class. They train five detectors: FasterRCNN [78], RetinaNet [55], CenterNet [104], FCOS [85], and the transformer-based DINO [98]. For the non-learned method, SAM is initially used to generate proposals, followed by employing DINO [11] and DINOv2 [68] to generate features for both proposal and template images, ultimately performing proposal matching and selection.
Synthetic-Real Test Sets. Li et al. [49] employ two benchmarks for evaluation. LineMod-Occlusion (LM-O) [5] includes 8 texture-less objects and 1,514 bounding boxes, primarily challenging due to significant occlusion. The YCB-Video (YCB-V) [8] features 21 objects and 4,125 bounding boxes, with pose variation as the main challenge. Since these datasets have real testing images without reference videos, Li et al. [49] generate synthetic videos using CAD models. We sample 16 synthetic template images per object from these videos.
RoboTools Benchmark. Li et al. [49] introduce RoboTools, a challenging real-world benchmark for robotics. It features 20 distinct instances, 9,109 annotations, and 24 complex scenarios. Compared to previous benchmarks [5, 8], RoboTools offers greater difficulty due to its cluttered backgrounds and significant pose variations. Additionally, the benchmark’s reference videos are made from real images and incorporate realistic lighting conditions, including shadows. 25 real template images of each instance are sampled from reference videos.
For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, Li et al. [49] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr. Initially, they generate open-world 2D proposals using their detection module [42]. For matching, different methods are employed to select the proposal with the highest score. In OLNDINO and OLNCLIP, they use robust features from pre-trained backbones [11, 75] and cosine similarity for matching. For OLNCorr, a matching head is designed based on correlation [57]. They also employ the class-level one-shot detectors OS2D [69] and BHRL [95]. To address the limitations of traditional 2D methods with pose variations and occlusions, Li et al. [49] introduced VoxDet which is based on a 3D voxel representation. To train these methods, Li et al. [49] developed the open-world instance detection (OWID-10k) dataset as a synthetic instance detection training set. It features 9,901 objects from ShapeNet [12] and ABO [19], each rendered into a 40-frame, 360-degree video using Blenderproc [22].
Evaluation Metrics. For detection, we evaluate our method with Average Precision (AP). AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05. This metric is the primary evaluation standard in the COCO Object Detection dataset [54]. Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
4.2 Segmentation Datasets
The BOP Challenge. We test our method using seven datasets from the BOP challenge [84]: LineMod Occlusion (LM-O) [5], T-LESS [35], TUD-L [36], IC-BIN [25], ITODD [26], HomebrewedDB (HB) [40], and YCB-Video [8]. The objects from these datasets vary widely. We use 42 template rendering images from CNOS [66], which are generated via BlenderProc [22].
Baselines. We compare our method with ZeroPose [13], CNOS [66], and SAM-6D [53]. These methods utilize proposals from SAM [47] or FastSAM [103] and use the $cls$ token from DINOv2 as the instance embedding for matching. In addition to adjusting the SAM hyperparameters to generate more proposals, SAM-6D enhances its performance by incorporating an appearance score, $s_{\text{appe}}$. Moreover, SAM-6D employs a Geometric Matching Score, $s_{\text{geo}}$, which considers the shapes and sizes of instances during matching by utilizing depth information.
Evaluation Metrics. For instance segmentation task, we evaluate our method using the Average Precision (AP) metrics. The AP is computed by averaging the precision scores at various IoU thresholds, from 0.50 to 0.95, increasing by 0.05 at each step.
Figure 4: Visual examples of our results using the Weight Adapter on the high-resolution and RoboTools datasets. Our approach detects specific object instances in cluttered scenes according to their real template images.
Table 2: Detection performance on the fully real dataset, RoboTools. Proposal indicates the object proposal model. GS stands for Grounded-SAM. OLN* is trained with the matching head while OLN employs fixed modules.
Method Proposal AP AP50 AP75
OS2D [69] N/A 2.9 6.5 2.0
DTOID [63] N/A 3.6 9.0 2.0
OLNCorr. [42, 57] OLN* 14.4 18.1 15.7
VoxDet [49] OLN* 18.7 23.6 20.5
NIDS-Net (Ours) GS 63.3 77.2 68.7
NIDS-Net + CA (Ours) GS 64.8 79.2 70.7
NIDS-Net + WA (Ours) GS 64.9 79.4 70.8
Table 3: Detection performance on the LM-O and YCB-V datasets. OLN* is trained alongside the matching head whereas OLN uses fixed modules. †means the model is trained on OWID and real images. ()($\cdot$)( ⋅ ) represents the number of instances.
LM-O (8) YCB-V (21) Average
Method Proposal Train AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 Time (s)
OS2D [69] N/A OWID 0.2 0.7 <0.1 5.2 18.3 1.9 2.7 9.5 1.0 0.189
DTOID [63] N/A OWID 9.8 28.9 3.7 16.3 48.8 42 13.1 38.9 4.0 0.357
OLNCLIP [42, 75] OLN OWID† 16.2 32.1 15.3 10.7 25.4 7.3 13.5 28.8 11.3 0.357
Gen6D [57] N/A OWID† 12.0 29.8 6.6 8.1 33.0 5.5 18.4 33.5 5.9 0.769
BHRL [95] N/A COCO 14.1 21.0 15.7 31.8 47.0 34.8 23.0 34.0 25.3 N/A
OLNCorr. [42, 57] OLN* OWID 22.3 34.4 24.7 24.8 41.1 26.1 23.6 37.8 25.4 0.182
OLNDino [42, 11] OLN OWID† 23.6 41.6 24.8 25.6 53.0 21.1 24.6 47.3 23.0 0.357
VoxDet [49] OLN* OWID 29.2 43.1 33.3 31.5 51.3 33.4 30.4 47.2 33.4 0.154
NIDS-Net (Ours) GS N/A 38.7 66.0 41.0 53.0 72.9 61.7 45.9 69.5 51.4 3.73
NIDS-Net + CA (Ours) GS N/A 39.2 67.0 41.4 53.9 74.1 62.7 46.6 70.6 52.1 3.71
NIDS-Net + WA (Ours) GS N/A 39.5 67.4 41.8 55.5 75.5 65.0 47.5 71.5 53.4 3.73
Table 4: Novel instance segmentation results on the seven core datasets of the BOP benchmark. We utilize Average Precision (AP) to compare these methods. ()($\cdot$)( ⋅ ) includes the number of instances. SAM6D (RGBD) exclusively utilizes RGB-D images, whereas other models employ only RGB images.
Method BOP Datasets
LM-O (8) T-LESS (30) TUD-L (3) IC-BIN (2) ITODD (28) HB (33) YCB-V (21) Mean
ZeroPose (SAM) [13] 34.4 32.7 41.4 25.1 22.4 47.8 51.9 36.5
CNOS (SAM) [66] 39.6 39.7 39.1 28.4 28.2 48.0 59.5 40.4
CNOS (FastSAM) [66] 39.7 37.4 48.0 27.0 25.4 51.1 59.9 41.2
SAM-6D (FastSAM) [53] 39.5 37.6 48.7 25.7 25.3 51.2 60.2 41.2
SAM-6D (FastSAM) + $s_{\text{appe}}$ [53] 40.6 39.3 50.1 27.7 29.0 52.2 60.6 42.8
SAM-6D (SAM) [53] 43.4 39.1 48.2 33.3 28.8 55.1 60.3 44.0
SAM-6D (SAM) +$s_{\text{appe}}$ [53] 44.4 40.8 49.8 34.5 30.0 55.7 59.5 45.0
SAM-6D (SAM, RGBD[53] 46.0 45.1 56.9 35.7 33.2 59.3 60.5 48.1
NIDS-Net (Ours) 42.9 43.0 52.0 30.5 28.8 56.6 59.7 44.8
NIDS-Net + CA (Ours) 42.8 47.5 50.0 29.7 27.6 62.7 63.5 46.3
NIDS-Net + WA(Ours) 43.2 49.2 53.5 30.5 29.4 62.4 64.3 47.5
NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 43.9 49.6 55.6 32.8 31.5 62.0 65.0 48.6
Table 5: Comparison of segmentation performance across various proposal models and embedding generation options. The first row is from CNOS [66].
Proposal Embedding BOP Datasets
LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
SAM cls token 39.6 39.7 39.1 28.4 28.2 48.0 59.5 40.4
SAM FFA 40.0 40.0 40.8 28.0 25.8 46.0 54.4 39.3
GS cls token 41.7 41.7 50.8 31.5 30 58 63 45.2
GS FFA 42.9 43.0 52.0 30.5 28.8 56.6 59.7 44.8
GS cls token + WA 42.5 48.2 50.7 31.3 32.0 60.5 62.8 46.9
GS FFA +WA 43.2 49.2 53.5 30.5 29.4 62.4 64.3 47.5
Figure 5: Comparison of segmentation results using CNOS, SAM6D, and our NIDS-Net on the YCB-V dataset. CNOS and SAM6D may misclassify some background regions or object parts as objects due to their reliance on SAM for proposal generation. NIDS-Net addresses this limitation with Grounded-SAM.
4.3 Benchmarking Results
Detection. The high-resolution dataset results are presented in Table 1. Our models dramatically outperform existing techniques, achieving the highest AP scores across all categories. Our basic method surpasses the top baseline by 17.7 AP. Additionally, as this dataset contains 100 instances with some similarities among them, our Weight Adapter boosts overall performance by 4.6 AP and improves the detection of small objects by 5.4 AP. For the RoboTools dataset, detailed in Table 2, our method outperforms the state-of-the-art VoxDet by 46.2 AP. Since RoboTools contains only 20 instances, our adapter has limited scope to enhance performance. Our approach achieves over 60 AP on these two fully real datasets.
For the Synthetic-Real datasets, LM-O and YCB-V, we present a comparison of our approach with other methods in Table 3. Our model outperforms VoxDet by 10.3 AP on LM-O and by 24.0 AP on YCB-V. The improved performance on YCB-V, which contains more instances, indicates that our adapter functions more effectively with increased instance variety.
Segmentation. CNOS [66] and SAM-6D [53] initially utilize SAM or FastSAM to obtain proposals and derive instance embeddings through the $cls$ token of Dinov2. In contrast, we generate proposals using Grounded-SAM and acquire embeddings through the FFA pipeline. The segmentation results for the seven principal datasets of the BOP challenge are detailed in Table 4. Overall, our method surpasses the state-of-the-art SAM-6D and achieves superior results compared to RGBD results.
Qualitative Result. We present detection results for both simple and complex scenes in Fig. 4. Fig. 5 displays a visual comparison of the segmentation results. Additional examples are available in the appendix.
Runtime. The tables of detection results include the runtime of our method. Grounded-SAM substantially decreases the number of proposals compared to SAM, thereby accelerating the entire process. Please see the appendix for segmentation runtime.
4.4 Ablation Study
Grounded SAM (GS) vs SAM. On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table 5. The results indicate that GS yields more precise bounding boxes and masks compared to using SAM alone. Furthermore, GS enhances efficiency by eliminating false object proposals, thereby reducing runtime.
FFA vs $cls$ token. To compare these two types of embedding generation, we present the results in Table 5. Our weight adapter enhances both two types of embeddings. Despite close segmentation results, FFA produces embeddings that possess greater adaptive potential, demonstrated by higher AP scores after adaption via our weight adapter. Additional ablation studies are provided in the appendix.
5 Discussions
In this study, we introduce NIDS-Net, a framework designed for novel instance detection and segmentation. We utilize Grounding DINO and SAM to generate precise foreground object proposals, as opposed to conventional naive region proposals. We also present the Weight Adapter, which refines features from a pre-trained vision model effectively and mitigates the risk of overfitting. With the adapter, template and proposal embeddings of different instances are separated to facilitate the subsequent matching. Our method surpasses other approaches significantly in detection performance and also excels in segmentation compared to existing methods.
However, there are limitations: given that our approach incorporates multiple pre-trained models, it requires greater computational resources compared to end-to-end detectors. Additionally, Grounding DINO may produce large bounding boxes that contain multiple objects, leading to detection failures. Moreover, our method tends to overlook heavily occluded objects with low confidence scores.
In this work, instances are represented by $K$ template embeddings. For future research, we will explore using a single, distinctive embedding for each instance that acts as its identifier to enable one-shot detection. This approach will allow the detector to identify and locate a target instance within a query image using just one template image. Additionally, developing a computationally efficient method for this process will be crucial for its application in robotics.
Acknowledgments
This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005.
Supplementary Material
6 Training Details
Detection. For detection datasets, the weight adapter is trained with a batch size of 1024, while the CLIP-Adapter is trained with a batch size of 512 to enhance performance. Both adapters are trained for the same number of epochs, as detailed in Table 6. To utilize Grounding DINO, a box threshold of 0.15 is set for the high-resolution and YCB-V datasets, and 0.10 for other datasets.
Table 6: The training epochs of different detection datasets.
High-resolution RoboTools LM-O YCB-V
Both Adapters 40 80 40 40
Segmentation. We combine all instances from seven core datasets of the BOP benchmark. We train both adapters with a batch size of 1344 (32 instances $\times$ 42 templates per instance) for 500 epochs. For Grounding DINO, a box threshold of 0.10 is set for all datasets.
7 More Ablation Study
FFA vs $cls$ token and Grounded SAM (GS) vs SAM in Detection Tasks. The detection capabilities of FFA and the $cls$ token are demonstrated using the high-resolution dataset, as presented in Table 9. For instance embedding generation, FFA achieves significantly greater enhancements compared to the $cls$ token. Furthermore, Table 9 illustrates that GS is effective in generating precise object proposals.
Table 7: Detection results using different ViT backbones of Dinov2. “reg” indicates DINOv2 with registers. The results are based on all testing images of the high-resolution dataset. WA Diff indicates the improvement attributed to the Weight Adapter.
Dinov2 backbone AP AP50 AP75 WA Diff
ViT-S/14 47.5 56.4 51.7 + 0.9 AP
ViT-S/14 + WA 48.4 57.5 52.9
ViT-B/14 54.7 65.2 60.1 + 3.2 AP
ViT-B/14 + WA 57.9 69.0 63.6
ViT-L/14 56.8 67.7 62.3 + 3.6 AP
ViT-L/14 + WA 60.4 71.8 66.4
ViT-L/14 reg 59.3 71.1 65.1 + 4.6 AP
ViT-L/14 reg +WA 63.9 76.6 70.6
Table 8: Detection results on the High-resolution real world detection dataset using various SAM variants.
SAM Variant AP AP50 AP75 Time (sec)
Mobile SAM (Tiny ViT) [97] 54.5 70.5 62.0 6.80
Mobile SAM (Tiny ViT) [97] + WA 58.2 75.0 66.5 6.73
HQ-SAM (ViT-H) [41] 59.8 71.5 65.7 7.38
HQ-SAM (ViT H) [41] + WA 63.8 76.2 70.5 7.31
SAM (ViT-H) (Ours) [47] 59.3 71.1 65.1 6.92
SAM (ViT-H) + WA (Ours) [47] 63.9 76.6 70.6 6.78
Table 9: The detection results on all images of the high-resolution dataset. “Proposal” refers to the object proposal method. “Embedding” denotes the method of instance embedding generation.
Proposal Embedding AP AP50 AP75
SAM $cls$ token 41.6 49.1 46.0
GS $cls$ token 54.9 65.4 60.1
GS FFA 59.3 71.1 65.1
Image encoder. Given the same object proposals with GS, we evaluate the FFA embeddings from different image encoders on the High-resolution dataset [81]. As illustrated in Table 10, DINOv2 exhibits superior performance attributable to its robust visual features. Fig. 6 presents the visual results of various image encoders.
Table 10: The instance embeddings of different image encoders.
Image Encoder Feature Dimension AP AP50 AP75
SAM (Vit-L) [47] 256 0.9 1.1 0.9
SAM (Vit-L) [47] + WA 256 0.7 0.8 0.7
DeiT III [86] (Vit-L) 784 2.9 3.6 3.1
DeiT III [86] (Vit-L) + WA 784 2.8 3.4 3.0
CLIP (Vit-L) [75] 1024 17.9 21.1 18.9
CLIP (Vit-L) [75] + WA 1024 20.4 24.0 21.8
DINOv2 (Vit-L) [68] 1024 56.8 67.7 62.3
DINOv2 (Vit-L) [68] + WA 1024 60.4 71.8 66.4
DINOv2 (Vit-L reg) [68] 1024 59.3 71.1 65.1
DINOv2 (Vit-L reg) [68] + WA 1024 63.9 76.6 70.6
Figure 6: Visual detection results on the High-resolution dataset using different image encoders.
Dinov2 backbones with adapter. Our Weight Adapter is compatible with various backbones of Dinov2. Notably, more powerful backbones, which offer a more effective feature space, enable our adapter to deliver greater improvements. Details of this comparison are provided in Table 7. Please refer to the Appendix for additional ablation studies.
Aggregation. For cls token embeddings, averaging the top $k$ highest (avg k) scores yields the best results [66, 53]. For FFA embeddings, the $\max$ aggregation function achieves optimal outcomes. The comparison of these aggregation functions is detailed in Table 11.
SAM variants. We evaluate different SAM variants on the High-resolution real world detection dataset. As illustrated in Table 8, SAM and our weight adapter achieves the superior performance. Given the precise bounding boxes, HQ-SAM[41] yields results comparable to those of SAM [47].
Table 11: Comparison of aggregation functions for segmentation performance. We report Average Precision (AP). avg k refers to averaging the top $k$ scores. All results are based on object proposals from Grounded SAM (GS).
Embedding Aggregation BOP Datasets
LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
$cls$ token avg k 41.7 41.7 50.8 31.5 30 58 63 45.2
$cls$ token $\max$ 42 37.4 45.9 30.2 28.9 54.9 61.5 43.0
FFA avg k 42.5 42 47.4 28.2 27.3 55.1 61.5 43.4
FFA $\max$ 42.9 43 52 30.5 28.8 56.6 59.7 44.8
Runtime. We compare the efficiency of existing methods for novel instance segmentation, as presented in Table 12. Our approach significantly reduces running time by proposing only high-quality bounding boxes.
Table 12: Runtime comparisons of various methods for novel instance segmentation.
Method Proposal Server Time (sec)
CNOS FastSAM Tesla V100 0.22
CNOS DeForce RTX 3090 0.23
SAM-6D (RGB) DeForce RTX 3090 0.25
SAM-6D (RGBD) DeForce RTX 3090 0.45
CNOS SAM Tesla V100 1.84
CNOS DeForce RTX 3090 2.35
SAM-6D (RGB) DeForce RTX 3090 2.28
SAM-6D (RGBD) DeForce RTX 3090 2.80
NIDS-Net (Ours) GS RTX A5000 0.49
NIDS-Net + CA (Ours) 0.48
NIDS-Net + WA (Ours) 0.48
NIDS-Net + WA + $s_{\text{appe}}$ (Ours) 0.48
8 Unseen Detection of BOP Benchmark
We compare our approach with ZeroPose [13], CNOS [66], and SAM-6D [53] for 2D unseen detection, as illustrated in Table 13. Our method outperforms the best RGB method by 2.5 AP and competes effectively with the top RGB-D method.
Table 13: Unseen instance detection results across the seven core datasets of the BOP benchmark, with all results reported as Average Precision (AP).
Method Proposal BOP Datasets
LM-O T-LESS TUD-L IC-BIN ITODD HB YCB-V Mean
ZeroPose [13] SAM 36.7 30.0 43.1 22.8 25.0 39.8 41.6 34.1
CNOS [66] SAM 39.5 33.0 36.8 20.7 31.3 42.3 49.0 36.1
CNOS [66] FastSAM 43.3 39.5 53.4 22.6 32.5 51.7 56.8 42.8
SAM-6D (RGB) [53] FastSAM 43.8 41.7 54.6 23.4 37.4 52.3 57.3 44.4
SAM-6D (RGBD) [53] SAM 46.6 43.7 53.7 26.1 39.3 53.1 51.9 44.9
SAM-6D (RGBD) [53] FastSAM 46.3 45.8 57.3 24.5 41.9 55.1 58.9 47.1
NIDS-Net (Ours) GS 44.9 42.8 43.4 24.4 34.9 54.8 56.5 43.1
NIDS-Net + WA(Ours) 44.9 48.9 46.0 24.5 36.0 59.4 62.4 46.0
NIDS-Net + WA (Ours) + $s_{\text{appe}}$ 45.7 49.3 48.6 25.7 37.9 58.7 62.1 46.9
9 More Qualitative Results
9.1 Adapter
To facilitate a comparison between the CLIP-Adapter and our Weight Adapter, we present a visual illustration in Figure 7. It is evident that the CLIP-Adapter alters the feature space and spoils the embeddings of background objects due to overfitting. In contrast, our Weight Adapter delivers robust embeddings within the original feature space.
9.2 Detection
We display the visual outcomes of our methodology employing the weight adapter on the LMO and YCB-V datasets in Figure 8. The gap between synthetic and real images results in some instances of detection failure. For example, some LM-O instances are not found with our method.
9.3 Segmentation
We present the visual outcomes of our approach using the weight adapter on the BoP datasets in Figures 9, 10, 11, and 12. These images demonstrate the effectiveness of our approach in cluttered environments. In some cases of T-LESS and IC-BIN datasets, Grounding DINO generates large bounding boxes which include multiple objects, causing under-segmentation. Furthermore, in IC-BIN and HB datasets, some heavily occluded objects with low confidence scores are overlooked by our method.
Figure 7: Comparison of different adapters on a hard scene image from the high-resolution dataset. Red arrows denote background objects that are erroneously classified as targets.
Figure 8: Visual detection results using the Weight Adapter on the LM-O and YCB-V datasets. Our approach detects object instances according to their synthetic template images.
Figure 9: Qualitative segmentation results on the LM-O and T-Less datasets.
Figure 10: Qualitative segmentation results on the TUD-L and IC-BIN datasets.
Figure 11: Qualitative segmentation results on the ITODD and HB datasets
Figure 12: Qualitative segmentation results on the YCB-V dataset.